# Beyond Numeric Awards: In-Context Dueling Bandits with LLM Agents

### v1

@@ 4.2 Experimental results
For brevity, we present our initial analysis focused on the Easy instance (Figure 2). The analysis is qualitatively similar for the Hard instance (Figure 8 in Appendix B.2). We use $\gamma=0.5$ for BTM, $f(K)=0.3K^{1.01}$ for RMED, $\eta=1$ for Self-Sparring, and $\alpha=0.51$ for RUCB, RCS and DTS. We analyze the results in terms of the strong and weak regret defined in Section 2.
Strong Regret. The strong regret results in Figure 2 (Left) reveal that among the LLMs, GPT-4 Turbo demonstrates the best performance in both the Easy and Hard instances. Comparing GPT-4 Turbo with the state-of-the-art baseline algorithms (DTS and Self-Sparring), we observe the following: (i) In the first $20$ time steps, GPT-4 Turbo exhibits similar or better performance than these baselines in both instances. (ii) Over the first $100$ time steps, GPT-4 Turbo maintains a performance level comparable to the state-of-the-art baselines in the Hard instance and slightly underperforms compared to the state-of-the-art baselines in the Easy instance, yet still outperforms all the other baseline algorithms. (iii) Throughout the entire time horizon, the strong regret of LLMs grows approximately linearly in both instances, indicating a non-converging behavior of the LLM agents. The results for the Hard instance (Figure 8) are qualitatively similar.
Weak Regret. In terms of the weak regret metric, Figure 2 (Right) indicates the exceptional performance of GPT-4 Turbo. It consistently outperforms all other LLMs and state-of-the-art baseline algorithms throughout the entire time horizon in both instances, exhibiting the lowest mean weak regret with small variance. The results for the Hard instance (Figure 8) are qualitatively similar.
To more thoroughly evaluate the capabilities of LLMs, we delve into the success and failure cases found in our experiments, with more results presented in Appendix B.2. In the following analyses, we focus on GPT-4 Turbo, which is the top-performing LLM as shown in Figures 2 and 8.
4.2.1 Advantages of LLM Agents
@@
Efficient Best Arm Identification. As evident in Figures 2 and 8 (Right), GPT-4 Turbo demonstrates strong zero-shot chain-of-thought (CoT) reasoning capabilities in terms of the weak regret. This can be attributed to its ability to quickly identify and consistently include the best arm in its duels. Figure 9 in Appendix B.2 illustrates the fraction of duels including the best arm across different time intervals. GPT-4 Turbo outperforms other LLMs and the state-of-the-art baselines throughout the entire timeline in both instances. These findings suggest that GPT-4 Turbo is able to effectively process the feedback obtained from duels and make informed decisions to identify and select the best arm, leading to superior weak regret performance.
Low Variance of Regret. GPT-4 Turbo demonstrates low variance compared to other LLMs and baseline algorithms. As shown in Figure 10 (Left) in Appendix B.2, GPT-4 Turbo exhibits the lowest average generalized variance of strong and weak regret in the Easy instance, suggesting that it can make consistent decisions when the preference structure is simple. For the Hard instance shown in Figure 10 (Right), GPT-4 Turbo has a variance level comparable to the state-of-the-art baseline algorithms (except BTM and SAVAGE, which are still in an early stage of exploration), indicating that its performance stability remains competitive even when faced with a more complex preference structure. This low variance result highlights GPT-4 Turbo’s ability to adapt to varying levels of difficulty and maintain a stable decision-making process in DB setting.
4.2.2 Disadvantages of LLM Agents
Figures 11 and 12 in Appendix B.2.2 present the arm selection traces of GPT-4 Turbo and state-of-the-art baselines on the Easy and Hard instances to analyze the failure cases. The trace plots suggest that GPT-4 Turbo exhibits a two-stage Explore-Then-Exploit pattern [4] with occasional robustness checks. Analysis of the two stages reveals two failure modes of LLMs:
@@ 4.2.2 Disadvantages of LLM Agents
Exploration Vulnerability. In the exploration stage, we observe that GPT-4 Turbo tends to quickly narrow down to a small subset of arms (usually containing the Condorcet winner) and repeatedly compare these arms. In contrast, the baselines exhibit more diverse and extended exploration patterns. This behavior suggests that GPT-4 Turbo may overestimate the quality of arms that win their initial comparisons based on limited historical data. Unlike the baselines, which have explicit exploration mechanisms, LLMs rely on their inherent randomness (via sampling from their output distribution) to explore. Based on these observations, we hypothesize that if GPT-4 Turbo happens to sample a sequence of comparisons that favors suboptimal arms early on, it can get stuck comparing these arms indefinitely. To test this hypothesis, we conducted experiments using noisy prompts with biased history. Our results in Figure 12 confirm that GPT-4 Turbo’s exploration strategy is indeed vulnerable to biased history initialization and can converge to local optima.
@@
Convergence Instability. Despite GPT-4 Turbo’s outstanding weak regret performance, it fails to consistently converge to a single best arm to duel against itself, even when the prompt setting explicitly calls for it. This behavior highlights a fundamental limitation of LLMs: they are primarily designed and trained for word token prediction rather than decision-making. Unlike baselines with explicit stopping conditions, GPT-4 Turbo relies on its inherent language modeling capabilities to determine when to stop exploring. Consequently, in the later exploitation stage, GPT-4 Turbo keeps comparing the same top arms without committing to a single winner (see Figure 11), reflecting an intrinsic uncertainty in its decision-making process for DB. The patterns learned during pre-training do not lead to converged decisions, even if the prompt includes emphatic commands, exclamation marks, quotation marks, or other forms of emphasis. This suggests that the language modeling objective alone may not be sufficient for LLMs to achieve optimal control in complex decision-making tasks like DB.
Type                                               Experimental Results Pros Best Arm Identification Condorcet Winner is identified and consistently selected
Results: Figure 28, and 9
Low Variance Low variance across instances of varying difficulty
Results: Figure 28, and 10
Cons Exploration Vulnerability Overestimates the quality of initially winning arms
Results: Figure 12
Convergence Instability Cannot guarantee convergence as a black-box language model
Results: Figure 28, and 11
Table 1: Summary of GPT-4 Turbo’s pros and cons in our experimental results.
Summary of Experimental Results. Table 1 summarizes the success and failure cases of using GPT-4 Turbo as an in-context decision-maker for the DB problem. GPT-4 Turbo can serve as an effective decision-maker in the short term, by quickly identifying and exploiting the best arm with low variance across different instances. However, the long-term performance is hindered by over-estimation bias in exploration stage and lack of convergence criterion in exploitation stage. This highlights the need for more robust and effective strategies to use LLMs.
@@
To further investigate the algorithmic behavior of LLMs in DB and develop more robust and effective strategies, we seek to answer the following questions:
@@
[Q1] Can we develop an LLM-augmented approach for DB with a theoretical guarantee?
@@ 5 Algorithm-Enhanced LLMs for Dueling Bandits
Classic DB algorithms such as Interleaved Filter 2 (IF2) [3] are known to be near-optimal, with matching regret upper and lower bounds up to multiplicative constants. To address the challenges identified in Table 1 of using LLM agents as standalone decision-makers for DB, we propose an LLM-augmented approach to demonstrate the possibility of taking advantage of both LLM agents and classic DB algorithms. Our algorithm, IF-Enhanced LLM, enjoys both a regret guarantee and strong empirical performance.
5.1 Algorithm Design of IF-Enhanced LLM
Figure 3: Main components of the proposed IF-Enhanced LLM agent in Algorithm 1 are illustrated: (i) The blue-colored part represents the LLM phase. (ii) The grey-colored part indicates the IF2 phase. (iii) The Match Arms Procedure and IF2 Procedure are detailed in Appendix A.1. (iv) The black arrows denote shared interactions between components. (v) The dotted arrows represent the input and output.
@@ 5.1 Algorithm Design of IF-Enhanced LLM
In this section, we present the design intuitions of IF-Enhanced LLM. We begin by discussing the limitations of a naive intervention approach and the desirable properties for an effective LLM augmentation framework. Based on these considerations, we identify IF2 [3] as the ideal candidate algorithm to serve as the foundation for IF-Enhanced LLM. Finally, we illustrate the components of IF-Enhanced LLM and provide a detailed algorithm description.
@@
Limitations of Naive Intervention. A straightforward approach to addressing the limitations of LLMs is to use a simple if-else condition that forces the LLMs to converge when they appear to select two identical arms, which we call the Convergence-Triggered (CT) intervention strategy. However, CT fails to guarantee the selection of the true Condorcet winner and can reinforce local optima (see Figure 13 in Appendix B.2 for a failure example). This suggests that relying on the LLM’s internal convergence behavior for such a mechanism is not robust, as the LLM’s exploration is largely driven by its inherent sampling noise rather than a structured exploration policy. Thus, handling this limitation with theoretical guarantees remains challenging.
@@
Desirable Properties for LLM Augmentation. To address [Q1], we seek an algorithmic framework with the following properties: (i) A clear, symbolic logical structure that allows for easy integration with LLM suggestions. (ii) A well-defined exploration-exploitation trade-off that leverages the LLMs’ exploration behavior while ensuring convergence. (iii) Strong theoretical guarantees to maintain robustness with various prompting scenarios.
IF2 as an Ideal Candidate. Among classic DB algorithms, IF2 stands out as particularly well-suited for LLM augmentation:
  • Its Explore-Then-Exploit [4] structure naturally aligns with the LLMs’ tendency to keep exploring without converging (see Figure 11), allowing for effective leveraging the LLMs’ exploration behavior while mitigating their exploration vulnerability and convergence instability (see Table 1).
  • Its symbolic representation of the algorithm’s logic enables clear integration of LLM suggestions at specific points without disrupting the overall structure and theoretical guarantees. In contrast, algorithms like Self-Sparring [7] are less symbolic, making them less suitable for direct LLM augmentation.
  • Its strong theoretical guarantees, with an expected regret bound of $O((K/\epsilon_{\mathrm{bad}})\log T)$ matching the DB problem’s lower bound of $\Omega((K/\epsilon_{\mathrm{bad}})\log T)$ up to constants (see Appendix A.2.1), and empirical performance (see Figures 2 and 8) provide a robust foundation, ensuring convergence and bounded regret.
By selecting IF2 as the base algorithm and leveraging its Explore-Then-Exploit pattern, we address [Q1]. IF2’s alignment with the observed behavior of LLMs (see Figure 11) allows for a natural integration that takes advantage of the LLMs’ exploration capabilities while maintaining theoretical guarantees.
@@
Algorithm Description. The procedures of the IF2-enhanced LLM agent are illustrated in Figure 3, presented in Algorithm 1 (see more details in Appendix A.1). IF-Enhanced LLM maintains a confidence parameter $\delta$ and a threshold parameter $\epsilon$ that control the algorithm’s confidence of matches between arms. The key components of IF-Enhanced LLM are as follows:
@@
  • Phase 1 (LLM Phase): Utilization of LLM recommended arms: The algorithm maintains a set of candidate arms $B$. Given two arms suggested by an LLM agent, the algorithm begins with finding a winner between them, denoted by $b_{\mathrm{LLM}}$. The winning arm $b_{\mathrm{LLM}}$ is then matched with each remaining arm $b\in B$, and $b$ is removed from $B$ if beaten by $b_{\mathrm{LLM}}$ with confidence $1-\delta$. This phase continues until $b_{\mathrm{LLM}}$ is defeated or all arms in $B$ have been matched. The variable TrustLLM is used to control the execution of the LLM phase, and it is set to False when $b_{\mathrm{LLM}}$ is defeated by another arm, indicating that the LLM’s suggestions are no longer trusted.
@@
  • Phase 2 (IF2 Phase): Roll back to IF2: If $b_{\mathrm{LLM}}$ is defeated, the algorithm switches to implementing one round of IF2 with an incumbent arm $b_{\text{{IF2}}}$ selected based on an estimated preference matrix $\hat{P}$.
@@
After Phase 2, the LLM-augmented agent repeats Phase 1 until $B$ only contains the best arm. Algorithm 1 and Figure 3 summarize the phases above, with detailed procedures defined in Appendix A.1.
Algorithm 1 LLM-augmented Agent: IF-Enhanced LLM
@@
$T,B=\{b_{1},\ldots,b_{K}\}$
@@
[…]
2        /* LLM Phase in Figure 3 (Lines 2 to 10) */ while TrustLLM do
             Prompt LLM to select $(b_{\mathrm{LLM}_{1}},b_{\mathrm{LLM}_{2}})$ from $B$ $(B,b_{\mathrm{LLM}})\leftarrow\textsc{Match Arms}(b_{\mathrm{LLM}_{1}},b_{\mathrm{LLM}_{2}})$ (Procedure 1)
3              /* Compare LLM selected arms */ for $b\in B$ do
                   $(B,b^{\prime})\leftarrow\textsc{Match Arms}(b_{\mathrm{LLM}},b)$ (Procedure 1)
[…]
@@
      $b_{\text{{IF2}}}\leftarrow\mathrm{argmax}_{b\in B}\min_{b^{\prime}\in B\setminus\{b\}}\hat{P}_{b,b^{\prime}}$
        /* Select the incumbent bandit */ Run IF2 (Procedure 2)
9        /* IF2 Phase in Figure 3 (Lines 11 to 12) */
@@
Output :  Estimated best arm $\hat{b}\in B$, Total number of comparisons $\hat{T}$
5.2 Theoretical Guarantees for IF-Enhanced LLM
In this section, we present two main theoretical results. First, we characterize the vulnerability of using standalone LLM agents for dueling bandits in Assumption 1 and Theorem 5.1. Then, we provide the theoretical guarantees of IF-Enhanced LLM in Theorem 5.2, demonstrating its robustness and efficacy.
Assumption 1 (Worst-Case Behavior).
Under the original prompt (see Figure 6), the worst-case behavior of an LLM agent in the dueling bandit setting is equivalent to a randomizer that selects actions uniformly at random.
Vulnerability of Standalone LLM Agents. Inspired by the adversarial corruptions framework introduced in [36] for the classic MAB problem, we investigate the vulnerability of standalone LLM agents in the DB setting under adversarial prompts. We consider an attacker with a budget $\Phi(T)$ who employs the following strategy: whenever the LLM agent selects the optimal arm $b^{*}$ for comparison, the attacker manipulates the input prompt to the LLM to eliminate $b^{*}$ from the duel with probability $p$ (where $0<p\leq 1$ is a constant), subject to the constraint of performing at most $\Phi(T)$ attacks over $T$ rounds. This adversarial strategy compels the LLM agent to select suboptimal arms, resulting in poor performance, as formalized in the following theorem with Assumption 1.

### v3

@@ 3.2 Experimental results
Figure 2: Comparisons between LLM agents and DB algorithms. Left and Right: strong and weak regret for the Transitive-Easy instance. Results for Transitive-Hard are in Figure 9.
@@
For brevity, we present our initial analysis focused on the Transitive-Easy instance (Figure 2). The analysis is qualitatively similar for the Transitive-Hard instance (Figure 9 in Appendix C.2). We use $\gamma=0.5$ for BTM, $f(K)=0.3K^{1.01}$ for RMED, $\eta=1$ for Self-Sparring, and $\alpha=0.51$ for RUCB, RCS and DTS. We analyze the results in terms of the strong and weak regret defined in Section 2. In the following sections, we will mainly focus on GPT-4 Turbo, which is our top-performing LLM, highlighting its success and failure modes.
@@
Emergence of in-context dueling bandits abilities. While GPT-3.5 Turbo and GPT-4 fail to solve the DB problem, GPT-4 Turbo consistently outperforms state-of-the-art DB baselines in weak regret on Transitive Case (see Figures 2 and 9). This reveals that the in-context dueling bandits abilities emerge as the general capabilities grow via standard training in general-purpose LLMs. Figure 13 (Left) illustrates the fraction of duels including the best arm across different time intervals. GPT-4 Turbo outperforms other LLMs and the DB baselines throughout the entire timeline. These findings suggest that GPT-4 Turbo can effectively process the preference feedback obtained from duels and make informed decisions to quickly identify and include the best arm in its duels.
Stable performance across different instances. GPT-4 Turbo demonstrates low variance compared to other LLMs and DB baselines across varying levels of difficulty. As shown in Figure 14, GPT-4 Turbo exhibits the lowest average generalized variance of strong and weak regret in both instances. This highlights its ability to maintain a stable decision-making process in DB.
Best Arm Identification: LLMs’ in-context dueling bandits abilities emerge as the general capabilities grow. The Condorcet Winner is consistently selected in duel via GPT-4 Turbo, leading to exceptional weak regret performance with minimal variance on Transitive Case.
@@
Exploration Vulnerability. In the exploration stage, we observe that GPT-4 Turbo tends to quickly narrow down to a small subset of arms (although usually containing the Condorcet Winner) and repeatedly compare these arms. In contrast, the baselines exhibit more diverse and extended exploration patterns. This behavior suggests that GPT-4 Turbo may overestimate the quality of arms that win their initial comparisons based on limited historical data. Unlike the baselines, which have explicit exploration mechanisms, LLMs rely on their inherent randomness (via sampling from their output distribution) to explore. Based on these observations, we hypothesize that if GPT-4 Turbo happens to sample a sequence of comparisons that favors suboptimal arms early on, it can get stuck comparing these arms indefinitely. To test this hypothesis, we conducted experiments using noisy prompts with biased history. Our results in Figure 16 confirm that GPT-4 Turbo’s exploration strategy is indeed vulnerable to biased history initialization and can converge to local optima.
@@
Exploitation Inability. Despite GPT-4 Turbo’s outstanding weak regret performance, it fails to consistently converge to a single best arm to duel against itself, even when the prompt setting explicitly calls for it. This behavior highlights a fundamental limitation of LLMs: they are primarily designed and trained for word token prediction rather than decision-making. Unlike baselines with explicit stopping conditions, GPT-4 Turbo relies on its inherent language modeling capabilities to determine when to stop exploring. Consequently, in the later exploitation stage, GPT-4 Turbo keeps comparing the same top arms without committing to a single winner (see Figure 3). This suggests that the language modeling objective alone may not be sufficient for LLMs to achieve optimal control in complex decision-making tasks like DB.
Lack of Robust Strategy: LLMs’ performance can be hindered by noisy and adversarial prompts due to overestimation bias in the exploration stage and the lack of convergence criteria in the exploitation stage.
Biased understanding of DB problem during pretraining. Our two best-performing LLMs, GPT-4 Turbo and o1-preview, exhibit systematic biases regarding the DB problem, likely due to a lack of exposure to similar tasks during pretraining. Specifically, they incorrectly assume that an arm cannot duel with itself (the convergence case), even when explicitly prompted to do so (see examples in Appendix C.1.3). This misunderstanding makes the DB problem as an out-of-distribution (OOD) task for LLMs, and in-context instructions fail to fully override this internal bias. Consequently, LLM agents cannot completely align with problem descriptions due to the inherent limitations of in-context learning, which cannot really generalize to OOD tasks (Wang et al., 2024b). Figure 13 supports these observations: o1-preview demonstrates better reasoning capabilities by transitioning from exploration to exploitation effectively and achieving lower strong regret than GPT-4 Turbo. However, its inference-time CoT mechanism reinforces its internal biased understanding of DB, resulting in bad weak regret performance due to the selection of two suboptimal arms in duels.
Systematic Biases: LLMs out-of-the-box lack a fundamental understanding of the DB problem and instead intuitively choose the next pair of arms to compare based on dueling history.
Scalability Limitation. To evaluate whether LLMs can generalize their exceptional weak regret performance, we conduct experiments from two perspectives: (i) Removing transitivity in preference structures: we change from transitive cases to intransitive cases that violate SST and STI (see Figures 10 and 11). The analysis of Intransitive-Easy and Intransitive-Hard is qualitatively similar: LLMs fail to replicate their long-term weak regret performance in transitive cases when faced with intransitive instances. However, their short-term weak regret performance (< 100 steps) remains exceptional; (ii) Increasing the number of arms: as illustrated in Figure 12, from $K=5$ to $K=10$, GPT-4 Turbo’s performance exhibits a noticeable long-term decline with the increase in $K$. While LLMs still beat all the DB baselines in the initial steps, they struggle to effectively infer the relative strengths among a larger number of arms in the long run. These findings suggest that while LLMs exhibit emergent abilities for relative decision-making rooted in linguistic knowledge, their effectiveness is only generalized to short-term scenarios. The long-term performance is hindered by larger number of arms and LLMs’ lack of fundamental understanding of intransitive cyclic structures.
To quantify this long-term scalability limitation and formally characterize the dueling bandit instances LLMs can handle, we introduce the concept of the Relative Decision Boundary (RDB). The RDB for a given LLM $m$ is defined as the set of problem difficulties $D$ for which the model achieves an acceptable level of weak regret across an effective short-term time horizon $T_{e}$, satisfying the condition:
$\text{RDB}(m)=\left\{(K,T,\Delta,T_{e})\ \Big{|}\ \mathsf{WR}(m,D(K,T,\Delta),T_{e})\leq\mathsf{R_{\text{th}}}\right\}.$ (2)
Here, $\mathsf{WR}(m,D,T_{e})$ represents the cumulative weak regret incurred by the model $m$ on a problem of difficulty $D$ over an effective short-term time horizon $T_{e}$, while $\mathsf{R_{\text{th}}}$ is a predefined threshold that quantifies acceptable weak regret performance. Overall, RDB is influenced by the inherent ability of $m$, the number of arms $K$, the transitivity $T$, the distinguishability between arms $\Delta$, and the effective short-term time horizon $T_{e}$.
Failure of long-term generalization: LLMs’ long-term strong and weak regret performance degrades when introducing intransitive preference structures or large number of arms. Success of short-term generalization: LLMs’ short-term weak regret performance remains surprisingly exceptional across all instances (see subfigures in 2,  9101112). We introduce Relative Decision Boundary (RDB) to describe the dueling bandit instances that LLMs can effectively handle.
Figure 3: Comparison of duel selection trajectories among GPT-4 Turbo, Self-Sparring, and DTS on the Transitive-Easy (Top Row) and Transitive-Hard (Bottom Row) instances. The decision trajectories of GPT-4 Turbo exhibit a clear pattern of continuous exploration without converging to the best arm. In contrast, Self-Sparring and DTS demonstrate structured exploration patterns and convergence properties.
As a summary, we found that in-context LLM agents’ linguistic prior allows them to quickly identify the Condorcet Winner from the dueling history in the short-term (for both Transitive Case and Intransitive Case), but it is vulnerable.
@@
To further investigate the algorithmic behavior of LLMs and develop more robust and effective in-context decision-making strategies, we seek to answer the following questions:
@@
[Q1] Can we develop an Algorithm-Enhanced in-context DB agent with a theoretical guarantee?
@@ 4 Algorithm-Enhanced LLMs for Dueling Bandits
Classic DB algorithms based on the Explore-then-Exploit framework, such as Interleaved Filter 2 (IF2) (Yue et al., 2012), are known to be near-optimal, with matching regret upper and lower bounds up to multiplicative constants. To address the challenges identified in Section 3.2 of using standalone LLM agents for DB, we propose an algorithm-enhanced approach: LLM with Enhanced Algorithmic Dueling (LEAD) to demonstrate the possibility of integrating off-the-shelf DB algorithms with LLM agents through fine-grained adaptive interplay. Our framework, LEAD, enjoys both a regret guarantee and strong empirical performance.
4.1 Algorithmic Design of LEAD
@@
In this section, we present the design intuitions of LEAD. We begin by discussing the limitations of a naive intervention approach and the desirable properties for an effective Algorithm-Enhanced LLM framework. Based on these considerations, we propose an agentic framework design LEAD, where we can incorporate any Explore-then-Exploit DB algorithms (Zoghi et al., 2014a). As an illustrative example, we use IF2 (Yue et al., 2012) to demonstrate how off-the-shelf algorithms can be integrated within LEAD and provide a detailed description.
@@
Limitations of Naive Intervention. A straightforward approach to addressing the convergence instability limitation of LLMs is to use a simple if-else condition that forces the LLMs to converge when they first exploit two identical arms, which we call the Convergence-Triggered (CT) intervention strategy. However, CT fails to guarantee the selection of the true Condorcet winner and can reinforce local optima (see Figure 17 in Appendix C.2 for a failure example). This suggests that relying on the LLMs’ internal convergence behavior to trigger the transition from exploration to exploitation is unreliable, as the LLMs are largely driven by its inherent sampling noise rather than a structured exploration policy. Thus, handling this limitation with theoretical guarantees remains challenging.
@@
Desirable Properties for LLM Augmentation. To address [Q1], we seek an algorithmic framework with the following properties: (i) A clear, symbolic logical structure that allows for easy integration with LLM & Algorithm suggestions; (ii) A well-defined exploration-exploitation trade-off that leverages the LLMs’ exploration behavior while ensuring convergence; (iii) Strong theoretical guarantees to maintain robustness with various prompting scenarios.
As a result, we find that the Explore-Then-Exploit structure is particularly well-suited for LLMs (see Appendix B.1 for a detailed illustration). By selecting an Explore-Then-Exploit DB algorithm as the foundation inside LEAD, we address [Q1]. As an example, we use IF2 as the base to illustrate the theoretical guarantee and empirical performance. This approach can be applied similarly to other algorithms with regret bounds in the Explore-Then-Exploit family.
Figure 4: Main components of the proposed LEAD agent in Algorithm 1 are illustrated: (i) The blue-colored part represents the LLM phase. (ii) The grey-colored part indicates the DB phase. (iii) The Algorithmic Procedures are detailed in Appendix B.2. (iv) The black arrows denote shared interactions between components. (v) The dotted arrows represent the input and output.
@@
Algorithmic Framework. The procedures of the LEAD are illustrated in Figure 4 and presented in Algorithm 1 (see more details in Appendix B.2). LEAD (IF2 base) maintains a confidence parameter $\delta$ and a threshold parameter $\epsilon$ that control the algorithm’s confidence of matches between arms. The key components of LEAD (IF2 base) are as follows:
@@
  • Phase 1 (LLM Phase): Utilization of LLM recommended arms: The agentic framework maintains a set of candidate arms $B$. Given two arms suggested by an LLM agent, the framework begins with finding a winner between them, denoted by $b_{\mathrm{LLM}}$. The winning arm $b_{\mathrm{LLM}}$ is then matched with each remaining arm $b\in B$. This phase continues until $b_{\mathrm{LLM}}$ is defeated or all arms in $B$ have been matched. The variable $\mathsf{TrustLLM}$ is used to control the execution of the LLM phase, and it is set to $\mathtt{False}$ when $b_{\mathrm{LLM}}$ is defeated by another arm, indicating that the LLM’s suggestions are no longer trusted.
@@
  • Phase 2 (IF2 Phase): Roll back to IF2: If $b_{\mathrm{LLM}}$ is defeated, the framework switches to implementing one round of IF2 with an incumbent arm $b_{\text{{IF2}}}$ selected based on an estimated preference matrix $\hat{P}$.
@@
After Phase 2, the algorithm-enhanced agent repeats Phase 1 until $B$ only contains the best arm. Algorithm 1 and Figure 4 summarize the phases above, with details delegated to Appendix B.2.
Algorithm 1 Algorithm-Enhanced LLM Agent: LEAD (IF2 base)
@@
Time horizon length $T$, arms $B=\{b_{1},\ldots,b_{K}\}$, incumbent arm $b_{\text{{IF2}}}$
@@
       $\mathsf{TrustLLM}\leftarrow\mathtt{True}$
2        /* LLM Phase in Figure 4 (Lines 2-10) */ while $\mathsf{TrustLLM}$ do
             Prompt LLM to select $(b_{\mathrm{LLM}_{1}},b_{\mathrm{LLM}_{2}})$ from $B$ $b_{\mathrm{LLM}}\leftarrow\textsc{Match Arms}(b_{\mathrm{LLM}_{1}},b_{\mathrm{LLM}_{2}})$ (Procedure 1)
3              /* Compare LLM arms */ for $b\in B$ do
                   $b^{\prime}\leftarrow\textsc{Match Arms}(b_{\mathrm{LLM}},b)$ (Procedure 1)
4                    /* Compare $b_{\mathrm{LLM}}$ with others */ if $b^{\prime}\neq b_{\mathrm{LLM}}$ then  $\mathsf{TrustLLM}\leftarrow\mathtt{False}$, continue
@@
      $\mathsf{StillTrust},B\leftarrow\textsc{Validate}(b^{\prime},B,\mathsf{TrustLLM})$ (Procedure 2) $b_{\text{{IF2}}},B\leftarrow$IF2$(b_{\text{{IF2}}},B)$ (Procedure 3)
9        /* IF2 Phase in Figure 4 (Lines 11-12) */
@@
11if $\mathsf{StillTrust}$ then return $b_{\mathrm{LLM}}$
12 else return $b_{\text{{IF2}}}$
4.2 Theoretical Guarantees for LEAD
In this section, we begin by characterizing the vulnerability of using standalone LLM agents for dueling bandits in Theorem 4.1. Then, we provide the theoretical guarantees of LEAD in Theorem 4.2 and  4.3, demonstrating its efficacy and convergence.